OpenStudy (anonymous):

Choose the equation that could be used to find three consecutive integers whose sum is 36. n + (n + 2) + (n + 4) = 36 n + (n + 1) + (n + 3) = 36 n + (n + 1) + (n + 2) = 36 n + (n − 1) + (n − 3) = 36

OpenStudy (zzr0ck3r):

what does consecutive mean?

OpenStudy (zzr0ck3r):

and which one of those looks like it has consecutive numbers in it?

OpenStudy (anonymous):

Consecutive integers are integers that follow each other in order

OpenStudy (anonymous):

but i do not know what to do thats the problem?

OpenStudy (zzr0ck3r):

1+2+3 = 1+(1+1)+(1+2) so if n = 1 1+2+3 = n+(n+1)+(n+2)